OpenStudy (anonymous):

I WILL MEDAL!! Maura had a pizza delivered to her home. The total amount she paid was $17.53. Included in the total amount were the menu price of the pizza, 6% sales tax on the price of the pizza, and a $3.75 delivery fee. What was the menu price of the pizza?

OpenStudy (lasttccasey):

Okay first you need to determine what exactly was taxed. Since it say the pizza was taxed, you can deduct the delivery fee first.

OpenStudy (anonymous):

it will be 14.78

OpenStudy (lasttccasey):

That is incorrect @Emma2002

OpenStudy (anonymous):

ok and

OpenStudy (lasttccasey):

$ 17.53 - 3.75 = T where T is the price which was taxed. From here, T = 1.06*P where P is the price on the menu. The 1.06 comes from 100% of the original price plus the 6% of tax = 106% or 1.06

OpenStudy (lasttccasey):

@help?!?!?! Do you understand?

OpenStudy (anonymous):

not really

OpenStudy (lasttccasey):

Okay let's take this backwards for a moment. Say you buy something at Wal-Mart for $10 (before tax). When you get to the register, you will have to pay tax of 6%. To find your total price, it will be the $10*(100 + 6)% = $10*1.06 = $10.60

OpenStudy (lasttccasey):

Since the delivery charge of the pizza isn't taxed, it's simply added to the subtotal price. So in the example I made up, if you were to say it costs $5 in gas to get home, you'd simply add that to the $10.60 to get $15.60. To solve your problem, we're simply doing this in reverse.

OpenStudy (lasttccasey):

Did that help any?

OpenStudy (anonymous):

yes

OpenStudy (lasttccasey):

Okay so from the equations above, what would be your price which is taxed, T?

OpenStudy (lasttccasey):

I have to go but to solve this real quick. 17.53 - 3.75 = 13.78 = T P = T/1.06 = 13 So the cost on the menu, before tax and delivery, is $13.